### v1

@@ 9 Conclusion
In this work, we have demonstrated a new technique $\mu$-GLANCE to search for microlensed GW signals from current detectors. In this method, we performed a model-independent search for microlensing signatures of the GWs by assuming a unlensed hypothesis and obtaining the best-fit GW strain fitting the data. The best-fit is used to calculate the residual signal in each detector. We cross-correlate the aligned residuals and obtain the residual cross-correlation. By comparing with the noise cross-correlation, we obtained a range of source and lensing parameters which would show up against the noise cross-correlation distribution. We also followed a lensing amplification model based approach for the joint exploration of the source and lensing parameters together. Notably, with the chosen simple lensing amplification model, we observe that GW source parameters are not degenerate with the lensing parameters. One important caveat of this technique is that, even if the model-dependent source and lens joint PE technique can work on single-detector events, the model-independent cross-correlation based-search cannot be performed using a single GW detector data.
@@
We would like to emphasize the necessity of using an accurate GW waveform model, that is able to capture all the essential features of compact binary (e.g. eccentricity, precession, higher modes). Otherwise, using inaccurate waveform model, we can observe non-zero residual cross-correlations even when there is no microlensing present in the GW data. To check whether the waveform model is producing residuals even when there is no microlensing, we can put it to a test by checking on a vast different BBH population. If the model is wrong, we may get residuals in many, if not all, of the GW signals and their residual cross-correlations will show up. This is how we can correct waveform model related systematics. However, since the inspiral phase is the longest duration part of the GW signal, and it contributes to most of the signal SNR, any waveform model that can capture the inspiral part well for a coalescing BBH event, can be considered as a good waveform model to search for the lensing signal. However, it is important to point out that though waveform mismodelling may lead to mismodelling of the microlensing parameters, but it is unlikely to impact the search microlensing search pipeline. This is because strain residual from best-fit will capture both waveform systematics and microlensing effects.
@@
Also, in this analysis, we have not incorporated non-stationarity in the strain noise [66, 67, 68]. The non-stationary noise behaviour at one detector is not correlated with the non-stationarity at a different detector, and the duration of a GW signal and the timescale over the which the noise PSD is changing are different. Therefore, we perform cross-correlation on such noise chunks from different detectors at a cross-correlation timescale ($\tau$) much shorter than the variation of the noise-PSD and much longer than a glitch. In this way, we can mitigate the noise fluctuations from generating significant deviations from zero residual cross-correlation  [^12] We also apply bandpass filter to remove very low frequencies where noise-fluctuation is very high and very high frequencies where the jitters are very faster than the signal oscillations .
@@
The current analytic lensing model we use in this analysis, captures most of the features present in microlensing amplification in general. However, we do not claim universality to this amplification template. It shows to be working nicely with the two test cases we used in 3. It may fail badly for some lens models present in the universe. However this would not impact the residual cross-correlation since, obtaining the best-fit strain at a detector and from there obtaining the residual and finally performing the residual cross-correlation – each step is done in a model independent approach. Therefore, when the amplification template fails, although hinders the joint parameter space exploration side, does not affect the microlensing detection side, which is a completely data-driven model-agnostic approach for a robust technique to detect microlensed GWs. However, we have not associated any of the microlensing parameters $a$, $b$, $f_{0}$, $k$ to physical space parameters in the lens system, like the offset parameter $y$, the mass of the lens $M_{l}$ or its density profile $\rho(\vec{r})$. This will be performed in a future work.
@@
With the gradual advancement of the ground-based GW detector sensitivities, and the inclusion of the next-generation like Cosmic Explorer (CE) [56], Einstein Telescope (ET) [57] or Laser Interferometer Space Antenna (LISA) [69], the microlensing signatures in the cross-correlation is expected to be predominant over noise cross-correlation distribution. This will allow microlensed GW signals to be detected at an unprecedented precision. The microlensing amplification can lead to inferring the intrinsic properties of the lensing object. Therefore, under the wave-optics regime, the microlensing aspects would allow us to probe the dark matter haloes and subhaloes, which are otherwise invisible to the electromagnetic signal. In this regard, $\mu$-GLANCE will play a pivotal role in both the detection and precise characterization of microlensing signatures, significantly enhancing our ability to identify and analyze these phenomena with greater efficiency.
Acknowledgments
@@
The authors are thankful to Sourabh Magare for reviewing the manuscript during the LSC Publications and Presentations procedure and providing useful comments. The authors express their gratitude to the ⟨data|theory⟩ Universe-Lab group-members for useful suggestions. This work is part of the ⟨data|theory⟩ Universe-Lab, supported by TIFR and the Department of Atomic Energy, Government of India. The authors express gratitude to the computer cluster of ⟨data|theory⟩ Universe-Lab for computing resources used in this analysis. We thank the LIGO-Virgo-KAGRA Scientific Collaboration for providing noise curves. LIGO, funded by the U.S. National Science Foundation (NSF), and Virgo, supported by the French CNRS, Italian INFN, and Dutch Nikhef, along with contributions from Polish and Hungarian institutes. The research leverages data and software from the Gravitational Wave Open Science Center, a service provided by LIGO Laboratory, the LIGO Scientific Collaboration, Virgo Collaboration, and KAGRA. Advanced LIGO’s construction and operation receive support from STFC of the UK, Max-Planck Society (MPS), and the State of Niedersachsen/Germany, with additional backing from the Australian Research Council. Virgo, affiliated with the European Gravitational Observatory (EGO), secures funding through contributions from various European institutions. Meanwhile, KAGRA’s construction and operation are funded by MEXT, JSPS, NRF, MSIT, AS, and MoST. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. We acknowledge the use of the following python packages in this work: NUMPY [70], SCIPY [71], MATPLOTLIB [72], GWSIM [73], ASTROPY[74], PYCBC [75], GWPY [76], LALSUITE [77], EMCEE [78] and CORNER [79], BILBY [80].
Notes:
The data files and result plots shown in this work will be available on the ⟨data|theory⟩ Universe-Lab GITHUB page. Its usage in a research work must be done with proper consent from the authors.
@@
Appendix A Appendix: Inclusion of amplification template phase term $\phi$ in the source and lens joint PE
@@
Figure 12: The figure shows the estimation of the amplification phase together with the microlensing parameters and source parameters as a part of joint parameter space exploration.
@@ Appendix B Appendix: Effect of Different Choices of $f_{0}$ Priors on Parameter Estimation
To understand whether we have effects on the posteriors caused by the choice of the priors, especially for the case when there was no microlensing injected in the data. We study the effect of the choice of the parameter $f_{0}$, for which we have two different priors: (i) $f_{0}=[20,1024]$ Hz and (ii) $f_{0}=[20,f_{\rm isco}]$ Hz, where $f_{\rm isco}$ is the innermost stable circular orbit frequency of the binary black holes and $1024Hz$ is the final frequency of the the frequency-series data.
We have plotted the frequency-domain data and the waveform model is constructed with the medians of the posteriors. To compare with we have shown in figure 13 that, we do not observe any direct effect of the choice of $f_{0}$ prior, since the medians of the microlensing perturbation parameter $b$ was very close to being zero. The medians of $b$ for both the cases with $f_{0}$ prior up to $1024$ Hz and $f_{\rm isco}$ are 0.0877 and 0.0764 respectively, therefore not showing any microlensing features prominently. We also obtained the joint distribution of the microlensing and source parameters when the signal was not microlensed, with those two different $f_{0}$ priors. The comparison of the choice of such different prior choices is shown in figure 14. Similarly, we obtained the joint distribution of the microlensing and source parameters when the signal was actually microlensed, with those two different $f_{0}$ priors. The comparison of the choice of such different prior choices is shown in figure 15.
Figure 13: Comparison between two different priors on $f_{0}$ on the reconstructed signal when the actual data had no microlensing features. We observe that the reconstructions do not show any significant differences in the microlensing features and therefore are very similar to each other.
Figure 14: The corner plot shows the joint distribution of source and microlensing parameters obtained for the two choices of priors mentioned, when the signal was not microlensed.
Figure 15: The corner plot shows the joint distribution of source and microlensing parameters obtained for the two choices of priors mentioned, when the signal was microlensed.
@@ Appendix C Appendix: False Alarm Rates Depending on Source and Lens Parameters
We fixed the oscillation parameter $f_{0}$ since from figure 9 we observed that the residual cross-correlation signal has no direct dependency on the value of $f_{0}$. Therefore, we only varied $a$, $b$, and $k$ to show at what FAR we can detect microlensing events.
Figure 16: The figure shows the variation of the FAR of microlensing signal detection depending on the signal strength and lensing parameters $a$, $b$, and $k$.
In figure 16, we varied the network matched-filtering SNR and observe the dependency of FAR of the microlensing detection, depending on the values of $a$, $b$, and $k$. We show on the first pane for $a=1.8$ at $b=0.3$ and $k=0.001$, it shows that close to network SNR of 70, we can have a FAR on at least one detector pair below $10^{-3}peryear$. For the case considering $a=1.2$, $b=0.3$ and $k=0.001$, the minimum network SNR is still at around 65 to detect a microlensing signal. This is because the effect of the strong lensing magnification $a$ has been absorbed in the luminosity distance while performing the parameter estimation for the best-fit. For the case with $a=1.2$, $b=0.4$ and $k=0.001$, we observe that the minimum network SNR is close to 45, where for the same $a$ and $b$ values, with $k=0.004$, it is moved back to 55. Finally, with $a=1.2$, $b=0.6$ and $k=0.001$, the minimum network SNR required is about 35.
From this study, we can conclude that microlensing is hard to detect with current generation GW detectors unless the SNR of the source is very high. With the inclusion of more detectors and better detector sensitivites such as LIGO-India [64], A# [81], CE[82] or ET[83] observatories, detection of microlensing events will be more feasible. The availability of more detectors and better detector sensitivity will help in obtaining cross-correlation signals between more pairs of detectors ($N_{\rm det}(N_{\rm det}-1)/2$, where $N_{\rm det}$ is the number of detectors) and will also help in reducing the FAR. As a result, will make it feasible for robust detection of the microlensing signatures from GW signals.

### v2

@@ 9 Conclusion
In this work, we have demonstrated a new technique $\mu$-GLANCE to search for WO microlensed GW signals from current detectors. In this method, we performed a model-independent search for WO microlensing signatures of the GWs by assuming a unlensed hypothesis and obtaining the best-fit GW strain fitting the data. The best-fit is used to calculate the residual signal in each detector. We cross-correlate the aligned residuals and obtain the residual cross-correlation. By comparing with the noise cross-correlation, we obtained a range of source and lensing parameters which would show up against the noise cross-correlation distribution. We also followed a lensing amplification model based approach for the joint exploration of the source and lensing template parameters together. Notably, with the chosen simple lensing amplification model, we observe that GW source parameters are not degenerate with the lensing parameters. One important caveat of this technique is that, even if the model-dependent source and lensing template parameters joint PE technique can work on single-detector events, the model-independent cross-correlation based-search cannot be performed using a single GW detector data.
@@
We would like to emphasize the necessity of using an accurate GW waveform model, that is able to capture all the essential features of compact binary (e.g. eccentricity, precession, higher modes). Otherwise, using inaccurate waveform model, we can observe non-zero residual cross-correlations even when there is no WO microlensing present in the GW data. To check whether the waveform model is producing residuals even when there is no WO microlensing, we can put it to a test by checking on a vast different BBH population. If the model is wrong, we may get residuals in many, if not all, of the GW signals and their residual cross-correlations will show up. This is how we can correct waveform model related systematics. However, since the inspiral phase is the longest duration part of the GW signal, and it contributes to most of the signal SNR, any waveform model that can capture the inspiral part well for a coalescing BBH event, can be considered as a good waveform model to search for the lensing signal. However, it is important to point out that though waveform mismodelling may lead to mismodelling of the WO microlensing parameters, but it is unlikely to impact the search WO microlensing search pipeline. This is because strain residual from best-fit will capture both waveform systematics and WO microlensing effects.
@@
Also, in this analysis, we have not incorporated non-stationarity in the strain noise (Coughlin, 2011; Mozzon et al., 2020, 2022). The non-stationary noise behaviour at one detector is not correlated with the non-stationarity at a different detector, and the duration of a GW signal and the timescale over the which the noise PSD is changing are different. Therefore, we perform cross-correlation on such noise chunks from different detectors at a cross-correlation timescale ($\tau$) much shorter than the variation of the noise-PSD and much longer than a glitch. In this way, we can mitigate the noise fluctuations from generating significant deviations from zero residual cross-correlation  [^13] We also apply bandpass filter to remove very low frequencies where noise-fluctuation is very high and very high frequencies where the jitters are very faster than the signal oscillations .
@@
The current analytic lensing model we use in this analysis, captures most of the features present in WO microlensing amplification in general. However, we do not claim universality to this amplification template. It shows to be working nicely with the two test cases we used in 3. It may fail badly for some lens models present in the universe. However this would not impact the residual cross-correlation since, obtaining the best-fit strain at a detector and from there obtaining the residual and finally performing the residual cross-correlation – each step is done in a model independent approach. Therefore, when the amplification template fails, although hinders the joint parameter space exploration side, does not affect the WO microlensing detection side, which is a completely data-driven model-agnostic approach for a robust technique to detect WO microlensed GWs. However, we have not associated any of the WO microlensing parameters $a$, $b$, $f_{0}$, $k$ to physical space parameters in the lens system, like the offset parameter $y$, the mass of the lens $M_{l}$ or its density profile $\rho(\vec{r})$. This will be performed in a future work.
@@
With the gradual advancement of the ground-based GW detector sensitivities, and the inclusion of the next-generation like Cosmic Explorer (CE) (Evans et al., 2021), Einstein Telescope (ET) (Maggiore et al., 2020) or Laser Interferometer Space Antenna (LISA) (Seoane et al., 2023), the WO microlensing signatures in the cross-correlation is expected to be predominant over noise cross-correlation distribution. This will allow WO microlensed GW signals to be detected at an unprecedented precision. The WO microlensing amplification can lead to inferring the intrinsic properties of the lensing object. Therefore, under the wave-optics regime, the WO microlensing aspects would allow us to probe the dark matter haloes and sub-haloes, which are otherwise invisible to the electromagnetic signal. In this regard, $\mu$-GLANCE will play a pivotal role in both the detection and precise characterization of WO microlensing signatures, significantly enhancing our ability to identify and analyse these phenomena with greater efficiency.
10 Acknowledgement
@@
The authors are thankful to Sourabh Magare for reviewing the manuscript during the LSC Publications and Presentations procedure and providing useful comments. The authors express their gratitude to the ⟨data|theory⟩ Universe-Lab group-members for useful suggestions. This work is part of the ⟨data|theory⟩ Universe-Lab, supported by TIFR and the Department of Atomic Energy, Government of India. The authors express gratitude to the computer cluster of ⟨data|theory⟩ Universe-Lab for computing resources used in this analysis. We thank the LIGO-Virgo-KAGRA Scientific Collaboration for providing noise curves. LIGO, funded by the U.S. National Science Foundation (NSF), and Virgo, supported by the French CNRS, Italian INFN, and Dutch Nikhef, along with contributions from Polish and Hungarian institutes. The research leverages data and software from the Gravitational Wave Open Science Center, a service provided by LIGO Laboratory, the LIGO Scientific Collaboration, Virgo Collaboration, and KAGRA. Advanced LIGO’s construction and operation receive support from STFC of the UK, Max-Planck Society (MPS), and the State of Niedersachsen/Germany, with additional backing from the Australian Research Council. Virgo, affiliated with the European Gravitational Observatory (EGO), secures funding through contributions from various European institutions. Meanwhile, KAGRA’s construction and operation are funded by MEXT, JSPS, NRF, MSIT, AS, and MoST. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. We acknowledge the use of the following python packages in this work: NUMPY (Harris et al., 2020), SCIPY (Virtanen et al., 2020), MATPLOTLIB (Hunter, 2007), ASTROPY(Collaboration et al., 2022), PYCBC (Nitz et al., 2024), GWPY (Macleod et al., 2021), LALSUITE (LIGO Scientific Collaboration et al., 2018), EMCEE (Foreman-Mackey et al., 2013) and CORNER (Foreman-Mackey, 2016), BILBY (Ashton et al., 2019b).
@@
Appendix A Appendix: Inclusion of amplification template phase term $\phi$ in the source and lensing effects joint Parameter Estimation
@@
Figure 12: The figure shows the estimation of the amplification phase together with the WO microlensing parameters and source parameters as a part of joint parameter space exploration.
@@ Appendix B Appendix: Effect of Different Choices of $f_{0}$ Priors on Parameter Estimation
To understand whether we have effects on the posteriors caused by the choice of the priors, especially for the case when there was no WO microlensing injected in the data. We study the effect of the choice of the parameter $f_{0}$, for which we have two different priors: (i) $f_{0}=[20,1024]$ Hz and (ii) $f_{0}=[20,f_{\rm isco}]$ Hz, where $f_{\rm isco}$ is the innermost stable circular orbit frequency of the binary black holes and $1024Hz$ is the final frequency of the the frequency-series data.
We have plotted the frequency-domain data and the waveform model is constructed with the medians of the posteriors. To compare with we have shown in figure 13 that, we do not observe any direct effect of the choice of $f_{0}$ prior, since the medians of the WO microlensing perturbation parameter $b$ was very close to being zero. The medians of $b$ for both the cases with $f_{0}$ prior up to $1024$ Hz and $f_{\rm isco}$ are 0.0877 and 0.0764 respectively, therefore not showing any WO microlensing features prominently. We also obtained the joint distribution of the WO microlensing template parameters and source parameters when the signal was not WO microlensed, with those two different $f_{0}$ priors. The comparison of the choice of such different prior choices is shown in figure 14. Similarly, we obtained the joint distribution of the WO microlensing template parameters and source parameters when the signal was actually WO microlensed, with those two different $f_{0}$ priors. The comparison of the choice of such different prior choices is shown in figure 15.
Figure 13: Comparison between two different priors on $f_{0}$ on the reconstructed signal when the actual data had no WO microlensing features. We observe that the reconstructions do not show any significant differences in the WO microlensing features and therefore are very similar to each other.
Figure 14: The corner plot shows the joint distribution of source and WO microlensing template parameters obtained for the two choices of priors mentioned, when the signal was not WO microlensed. The parameter space occupied by these two cases, although very similar for the source parameters, shows slight disagreement for the lensing parameters.
Figure 15: The corner plot shows the joint distribution of source and WO microlensing parameters obtained for the two choices of priors mentioned, when the signal was WO microlensed. We observe that the choice of prior does not shift any of source or lensing parameters, if not shifted marginally.
@@ Appendix C Appendix: False Alarm Rates Depending on Source and Lens Parameters
We fixed the oscillation parameter $f_{0}$ since from figure 9 we observed that the residual cross-correlation signal has no direct dependency on the value of $f_{0}$. Therefore, we only varied $a$, $b$, and $k$ to show at what FAR we can detect WO microlensing events.
Figure 16: The figure shows the variation of the FAR of WO microlensing signal detection for a couple of more variations on lensing amplification parameters $a$, $b$, and $k$. This helps us understand for a wide variety of lensing amplification effects, how confidently we are going to claim a lensing detection for a given network SNR.
In figure 16, we varied the network matched-filtering SNR and observe the dependency of FAR of the WO microlensing detection, depending on the values of $a$, $b$, and $k$. We show on the first pane for $a=1.8$ at $b=0.3$ and $k=0.001$, it shows that close to network SNR of 70, we can have a FAR on at least one detector pair below $10^{-3}peryear$. For the case considering $a=1.2$, $b=0.3$ and $k=0.001$, the minimum network SNR is still at around 65 to detect a WO microlensed signal. This is because the effect of the strong lensing magnification $a$ has been absorbed in the luminosity distance while performing the parameter estimation for the best-fit. For the case with $a=1.2$, $b=0.4$ and $k=0.001$, we observe that the minimum network SNR is close to 45, where for the same $a$ and $b$ values, with $k=0.004$, it is moved back to 55. Finally, with $a=1.2$, $b=0.6$ and $k=0.001$, the minimum network SNR required is about 35.
From this study, we can conclude that WO microlensing is hard to detect with current generation GW detectors unless the SNR of the source is very high. With the inclusion of more detectors and better detector sensitivities such as LIGO-India (Iyer et al., 2011), A# (Peter Fritschel et al., 2024), CE(Hall et al., 2021) or ET(Punturo et al., 2010) observatories, detection of WO microlensed events will be more feasible. The availability of more detectors with better detector sensitivity will help in obtaining cross-correlation signals between more pairs of detectors ($N_{\rm det}(N_{\rm det}-1)/2$, where $N_{\rm det}$ is the number of detectors) and will also help in reducing the FAR. As a result, it enables for a robust detection of the WO microlensing signatures from GW signals.